OpenStudy (anonymous):

Find the magnitude and directional angle for the following vector. <-4,4sqrt3>

OpenStudy (anonymous):

 then it would have angle using trig functions, for example, \[\tan^{-1} \theta = 4\sqrt{3} \div -4\] Then your magnitude would be determined by \[|v| = \sqrt{ x ^{2} + y ^{2}} \rightarrow |v| =\sqrt{ 4^2 + (4\sqrt{3})^2}\]